Do It! Review 18-2 Westerville Company accumulates the following data concerning a mixed cost, using units produced as the activity level March April May June July Units Produced 9,950 9,050 10,400 8,900 9,550 Total Cost $13,240 12,690 13,940 12,290 13,140 Compute the variable cost per unit using the high-low method. (Round variable cost per unit to 2 decimal places e.g. 12.25.) Variable cost per unit LINK TO TEXT Compute the fixed cost elements using the high-low method. (Round answer to 0 decimal places, e.g. 5,275.) Fixed cost LINK TO TEXT Estimate the total cost if the company produces 8,000 units. (Round answer to 0 decimal places, e.g. 5,275.) Total cost to produce 8,000 units
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Cost at highest level of activity          13,940.00
Cost at Lowest level of activity          12,290.00
Highest level of activity          10,400.00
Lowest Level of activity            8,900.00
Variable cost per unit = (13940 - 12290)/(10400 - 8900)
Variable cost per unit = 1650/1500
Variable cost per unit = $1.10
Fixed Costs = 13940 - 10400*1.10
Fixed Costs = 13940 - 11,440
Fixed Costs = 2500
Total cost to produce 8000 Units = 8000*1.10 + 2500
Total cost to produce 8000 Units = 8800 + 2500
Total cost to produce 8000 Units = 11,300